# Rethinking Creativity Evaluation: A Critical Analysis of Existing Creativity Evaluations

### v2

@@ 5.2.1 Creativity index: Idea originality $\neq$ unique n-grams in the content
crossley2016idea explore the relationship between linguistic features and idea generation, operating under the assumption that "essays with greater originality would contain a greater number of unique n-grams," where originality is defined as "how different an individual’s ideas are from others’ ideas." However, our results, as reflected by CI in Problem-solving and Research ideation, show that texts containing original ideas do not necessarily exhibit a high frequency of unique n-grams. On the other hand, merrill2024evaluating assess the novelty of LLM-generated versus human-written text through n-gram comparisons, where "novelty" is defined as the presence of n-grams not found in a reference corpus. While this approach aligns with our findings in Creative writing, where LLM-generated text contains fewer unique n-grams, it is important to note the presence of unique n-grams primarily indicates lexical rarity or uncommon phrasing, rather than genuine novelty or creativity at the conceptual or ideational level.
@@ 5.2.2 Creativity index: What affects the credibility of the creativity index?
Range of L-uniqueness. In the original study, the choice of L (the range of n-gram lengths) used for CI computation is condition-specific. As shown in Figure 2, CI is particularly sensitive to the choice of L. When the range is expanded from 5–7 to 5–11, the CI score differences between human-written plots and LLM-generated content become negligible as shown in Figure 3. This sensitivity limits CI’s generalizability across different domains and content types and raises concerns about potential bias in its ability to reflect true linguistic creativity.
Figure 2: L-uniqueness v.s. minimum n-gram length L. Creativity index is sensitive to L.
Figure 3: CI value v.s. L-uniqueness Range. As L-uniqueness increases, the CI difference decreases.
5.2.3 Syntactic templates: differences in LLM-generated v.s. human-written content
@@
Figure 4: Template rate by n-gram Template. As the top 100 most common templates are selected, their actual occurrence rate varies with n.
@@ 5.2.4 LLM-as-a-judge’s limitation in reflecting human-perceived creativity
Figure 6: LLM evaluation results. (a) Problem-solving: LLMs tend to rate the solutions to be conventional; (b) Creative writing: LLMs tend to rate the plots to be non-original.
@@ 5.2.5 Improving LLM-as-a-judge via chain-of-thought and rubric-based prompting
Chain-of-thought (CoT; wei2022chain) assists LLMs in better adhering to evaluation instructions and consequently refining their final output scores. For example, in the third round of dialogue, the Claude agent revises (soundness, presentation, contribution, overall) evaluation scores from (3.5, 3, 3.5, 7) to (3, 3, 3, 6), with an explanation "…However, for a conference with only 30% acceptance rate, I must be stringent. The paper, while valuable, doesn’t appear to represent a breakthrough…, warranting a "Weak Accept" rating." This demonstrates the model’s ability to self-correct and adjust its assessment based on the provided context and constraints.
@@ 6 Conclusion
Future research can build on our findings by collecting high-quality, orthogonal datasets in underexplored domains such as musical composition, humor, or product design. These datasets would complement our current benchmarks and support the development of generalizable, multi-domain, conceptually aware creativity judges that capture signals beyond surface-level patterns.
@@ Appendix A Data curation for scientific ideation
Alongside the accept or reject status, we collect the presentation, soundness, contribution, and overall scores for each paper. Scientific contribution is commonly used to represent creativity in the context of research ideation (zhao2025review). To keep the focus on the research framing and ideation components, we filter for the Introduction section of papers, select papers based on the average of their overall review scores, and utilize soundness and presentation scores as control variables to reduce the influence of other factors.
@@ Appendix B CI implementation
For the results in Table 2, we compute CI with $L$-uniqueness in the range $5\leq L\leq 7$, following Appendix B.3 of lu2025salieri. We observe the same saturation behavior for $L\geq 7$ (values $\approx$ 1), and therefore adopt the same setting. We also consider the near-verbatim match method using Word Mover’s Distance. However, this approach has computational complexity $O(|d|^{2}|w|)$ for each document $d$ and n-gram $w$, making it impractical for our large, diverse datasets. Moreover, (lu2025salieri) reported similar human-vs-LLM gaps whether using verbatim or verbatim+semantic matching. Therefore, for consistency and scalability, we use the verbatim method for our main experiments.
@@ Appendix C Label verification for creative writing
In the dataset from prior work (tian2024a), turning points (markers of a story’s pacing) were shown to have highly significant differences at tp4 (mean_diff = 3.59-3.59, $p=8.7\times 10^{-8}$) and tp5 (mean_diff = 4.19-4.19, $p=5.4\times 10^{-8}$), indicating that LLM narratives reach these late-stage beats much earlier than human ones. Similarly, the distribution of story arcs (general plotlines such as “Man in Hole” or “Rags to Riches”) differs dramatically ($p=1.9\times 10^{-11}$). To further support our assumption, we have run some additional pairwise human annotations.
@@
Appendix D Human annotation details and instructions
[…]
@@ Appendix D Human annotation details and instructions
Research ideation: "Below are some paper introductions. Select the one with higher contribution, bringing more new scientific breakthoughs and innovation to the research community."
@@ Appendix E Metrics yield misaligned performance across the same concept
SOURCE Exact CI↑ PPL↑ Syntactic Templates↓ LLM-as-a-Judge↑
Mean std. Mean std. CR-POS↓ $\geq$1 Template (%)↓ Mean std.
@@ F.1 Prompt selection
Roleplaying with constraints v.s. Multi-aspect rubric-based prompt: Simply instructing the model does not result in more discriminative judgments. For example, in scientific ideation, contribution scores under constraints are indistinguishable from those produced with score-based novelty prompts. In contrast, multi-aspect rubric with orthogonal dimensions better isolates confounding factors such as presentation and soundness, and provides judgments more specifically targeted at scientific creativity, as discussed in Section 5.2.5.
@@ F.4 Potential data contamination
Group GT Label Human Scores Claude 3 Haiku GPT-o1 GPT-4o GPT-o3-mini Claude Sonnet 3.7
@@
Table 5: Performance of LLM judges with different cutoffs. Comparison of human and models’ evaluation across groups (red = groups misaligned with ground-truth (GT) labels; green = groups aligned with GT labels). Although Claude Sonnet 3.7 achieves the strongest alignment, this may be due to potential training data contamination rather than a genuine ability to evaluate conceptual novelty.
@@ Appendix G LLM evaluation prompts
Domain - Creative writing f”’Given a movie plot: “‘{entry[’Content’]}“‘ Analyze the plot in terms of the following four aspects: Background Setup, Logic, Development, and Originality. For each aspect, consider the following questions: Background Setup: How common or unique is the background setting? Logic: Is the timeline clear and are the cause-and-effect relationships logical? Development: Are the twists well-developed? Is the buildup to the climax sufficient and engaging? Originality: How predictable is the plot? How original are the events and developments? For each aspect, discuss the strengths and weaknesses of the plot. Provide a thoughtful and critical analysis.”’, ”’Given the movie plot and your analyzation across four aspects, which The criteria for each aspect are as follows: A. Background Setup 1. Very common or clichéd background, lacks uniqueness. 2. Somewhat common, but with a few unique elements. 3. Generally unique, with some familiar elements. 4. Mostly unique background and setup. 5. Exceptionally unique background setup. B. Logic 1: Poor logic, many inconsistencies or plot holes. 2: Somewhat confusing or unrealistic transitions. 3: Generally coherent with minor issues. 4: Mostly logical and well-structured. 5: Strong internal logic and smooth cause-effect flow. C. Development 1: Underdeveloped or abrupt changes. 2: Weak buildup, limited depth. 3: Average development, a few effective moments. 4: Good buildup with engaging twists. 5: Excellent pacing, buildup, and well-executed climax. D. Originality 1: The plot is highly predictable or derivative. 2: Mostly familiar with few surprises. 3: Balanced mix of familiar and fresh ideas. 4: Refreshingly different with clever elements 5: Exceptionally creative and surprising or inspiring throughout. Ensure that each aspect is orthogonal to the others—for example, the uniqueness of background setup should not affect the evaluation of plot originality and development. You are a rigorous reviewer. Provide the scores in the following format: [[Background Setup: X]], [[Logic: X]], [[Development: X]], [[Originality: X]].”’
@@
Domain - Problem-solving f”’Given the problem: “‘{entry[’Problem’]}“‘ and the solution: “‘{entry[’Content’]}“‘, analyze the following three aspects, taking into account the corresponding dimensions: 1. Problem Difficulty: Is the problem conceptually complex, or does it involve multiple steps or layers of reasoning? Are there hidden constraints or non-obvious elements that make the problem harder? 2. Unconventionality of the Solution: Does the solution follow a standard method, or does it use a surprising or creative strategy? Would this solution surprise someone familiar with typical methods for this problem? Are the tools, concepts, or constraints used in unexpected ways — beyond their typical or intended function? 3. Efficiency of the Solution: Does it reach the goal using the fewest steps or resources possible? Could the solution be improved or simplified without losing correctness? Carefully summarize and provide a thorough assessment of each aspect.”’, f”’Given the problem, solution, and your analyzation across three aspects, provide a score for each aspect accordingly, which The criteria for each aspect are as follows: A. Problem Difficulty 1: Very Easy (trivial or commonly seen problems; requires little reasoning) 2: Easy (straightforward with minimal abstraction or complexity) 3: Moderate (requires some thought or mild creativity; manageable) 4: Challenging (requires significant thought, creative insight, or multiple reasoning steps) 5: Very Challenging (highly complex or abstract; involves deep insight or several layers of reasoning) B. Unconventionality 1: Very Conventional (the most expected approach; standard method anyone would think of first) 2: Conventional (follows a known method that is not surprising, but slightly less obvious) 3: Moderately Unconventional (mix of familiar and unexpected methods or framings) 4: Unconventional (surprising to an expert; includes a creative or less typical approach) 5: Highly Unconventional (reframes the problem or uses tools and constraints in a deeply creative, unexpected way) C. Efficiency 1: Very Inefficient (excessive or redundant steps; far from optimal) 2: Inefficient (works, but could clearly be improved with fewer steps or resources) 3: Moderately Efficient (reasonable efficiency; could still be optimized) 4: Efficient (solid balance of steps and resources; minimal waste) 5: Very Efficient (achieves the goal in the fewest possible steps or resources with no clear room for improvement) Ensure that each aspect is orthogonal to the others—for example, Unconventionality should not be affected by the Efficiency or the Problem Difficulty. Strictly follow the criteria and provide the scores in the following format: [[Problem Difficulty: X]], [[Unconventionality: X]], [[Efficiency: X]].”’
@@
Domain - Research Ideation f”’ You are a reviewer for a top conference, the acceptance rate for this conference is only around 20-30%. Given the introduction of the paper: “‘entry[’Content’]“‘, carefully summarize and provide a thorough assessment of the strengths and weaknesses of the paper, touching on each of the following dimensions: 1. Originality: Are the tasks or methods new? Is the work a novel combination of well-known techniques? Is it clear how this work differs from previous contributions? 2. Quality: Is the submission technically sound? Are claims well supported (e.g., by theoretical analysis or experimental results)? Are the methods used appropriate? Is this a complete piece of work or work in progress? Are the authors careful and honest about evaluating both the strengths and weaknesses of their work? 3. Clarity: Is the submission clearly written? Is it well organized? Does it adequately inform the reader? 4. Significance: Are the results important? Are others (researchers or practitioners) likely to use the ideas or build on them? Does the submission address a difficult task in a better way than previous work? Does it advance the state of the art in a demonstrable way? Does it provide unique data, unique conclusions about existing data, or a unique theoretical or experimental approach? 5. Limitations: Have the authors adequately addressed the limitations and potential negative societal impact of their work? If not, please include constructive suggestions for improvement.”’, ”’Given the introduction of the paper and the strengths and weaknesses you summarized, analyze the four aspects of the paper repectively: Soundness, Presentation, Contribution, and Overall Assessment. The criteria for each aspect are as follows: A. Soundness: the soundness of the technical claims, experimental and research methodology and on whether the central claims of the paper are adequately supported with evidence. 4 excellent; 3.5; 3 good; 2.5; 2 fair; 1.5; 1 poor. B.Presentation: the quality of the presentation. This should take into account the writing style and clarity, as well as contextualization relative to prior work. 4 excellent; 3 good; 2 fair; 1 poor. C. Contribution: the quality of the overall contribution this paper makes to the research area being studied. Are the questions being asked important? Does the paper bring a significant originality of ideas and/or execution? Are the results valuable to share with the broader research community. 4 excellent; 3 good; 2 fair; 1 poor. D. Overall: 10: Award quality: Technically flawless paper with groundbreaking impact on one or more areas, with exceptionally strong evaluation, reproducibility, and resources, and no unaddressed ethical considerations. 9: Very Strong Accept: Technically flawless paper with groundbreaking impact on at least one area and excellent impact on multiple areas, with flawless evaluation, resources, and reproducibility, and no unaddressed ethical considerations. 8: Strong Accept: Technically strong paper with, with novel ideas, excellent impact on at least one area or high-to-excellent impact on multiple areas, with excellent evaluation, resources, and reproducibility, and no unaddressed ethical considerations. 7: Accept: Technically solid paper, with high impact on at least one sub-area or moderate-to-high impact on more than one area, with good-to-excellent evaluation, resources, reproducibility, and no unaddressed ethical considerations. 6: Weak Accept: Technically solid, moderate-to-high impact paper, with no major concerns with respect to evaluation, resources, reproducibility, ethical considerations. 5: Borderline accept: Technically solid paper where reasons to accept outweigh reasons to reject, e.g., limited evaluation. Please use sparingly. 4: Borderline reject: Technically solid paper where reasons to reject, e.g., limited evaluation, outweigh reasons to accept, e.g., good evaluation. Please use sparingly. 3: Reject: For instance, a paper with technical flaws, weak evaluation, inadequate reproducibility and incompletely addressed ethical considerations. 2: Strong Reject: For instance, a paper with major technical flaws, and/or poor evaluation, limited impact, poor reproducibility and mostly unaddressed ethical considerations. 1: Very Strong Reject: For instance, a paper with trivial results or unaddressed ethical considerations.”’, ”’Given the introduction of the paper, the strengths and weaknesses you summarized, and your analysis of the four aspects, provide a score for each aspect accordingly. Ensure that each aspect is orthogonal to the others—for example, Presentation should not affect the evaluation of Soundness or Contribution. You are a rigorous reviewer, and the conference has an acceptance rate of no more than 30%. Provide the scores in the following format:[[Soundness_llm: X]], [[Presentation_llm: X]], [[Contribution_llm: X]], [[Overall_llm: X]].”’

### v3

@@ 5.2.1 Creativity Index: Idea originality $\neq$ unique n-grams in the content
Crossley et al. (2016) explore the relationship between linguistic features and idea generation, operating under the assumption that "essays with greater originality would contain a greater number of unique n-grams," where originality is defined as "how different an individual’s ideas are from others’ ideas." However, our results, as reflected by CI in Problem-solving and Research ideation, show that texts containing original ideas do not necessarily exhibit a high frequency of unique n-grams. On the other hand, Merrill et al. (2024) assess the novelty of LLM-generated versus human-written text through n-gram comparisons, where "novelty" is defined as the presence of n-grams not found in a reference corpus. While this approach aligns with our findings in Creative writing, where LLM-generated text contains fewer unique n-grams, it is important to note the presence of unique n-grams primarily indicates lexical rarity or uncommon phrasing, rather than genuine novelty or creativity at the conceptual or ideational level.
@@ 5.2.2 Creativity Index: What affects the credibility of the Creativity Index?
Range of L-uniqueness. In the original study, the choice of L (the range of n-gram lengths) used for CI computation is condition-specific. As shown in Figure 2, CI is particularly sensitive to the choice of L. When the range is expanded from 5–7 to 5–11, the CI score differences between human-written plots and LLM-generated content become negligible as shown in Figure 3. Table 3 demonstrates that this sensitivity also persists across different reference corpora, specifically (1) Pile-val and (2) Dolma-v1.6-sample, where CI scores are computed under multiple L-ranges (5–5, 5–7, 5–9, and 5–11). This sensitivity limits CI’s generalizability across different domains and content types and raises concerns about the discriminative power and potential bias in its ability to reflect true linguistic creativity.
(1) Corpus: Pile-val
Domain Truth label 5–5 5–7 5–9 5–11 Creative writing human expert 0.65 0.82 0.89 0.92 LLM-generated 0.47 0.71 0.82 0.87 Problem-solving unconventional 0.69 0.86 0.91 0.94 conventional 0.69 0.85 0.91 0.94 Research ideation accepted 0.65 0.81 0.88 0.91 rejected 0.64 0.80 0.87 0.91
(2) Corpus: Dolma-v1.6-sample
Domain Truth label 5–5 5–7 5–9 5–11 Creative writing human expert 0.48 0.72 0.82 0.87 LLM-generated 0.26 0.53 0.69 0.78 Problem-solving unconventional 0.52 0.76 0.85 0.89 conventional 0.52 0.75 0.85 0.89 Research ideation accepted 0.50 0.71 0.81 0.86 rejected 0.50 0.71 0.81 0.86
Table 3: Performance across corpora and domains. Scores increase with wider ranges, with substantial overlap between labels across domains.
Figure 2: L-uniqueness vs. minimum n-gram length L. Creativity Index is sensitive to L.
Figure 3: CI value vs. L-uniqueness range. As L-uniqueness increases, the CI difference decreases.
5.2.3 Syntactic templates: Differences in LLM-generated vs. human-written content
@@
Figure 4: Template rate by n-gram tunemplate. As the top 100 most common templates are selected, their actual occurrence rate varies with n. Note that the sudden spike at n=8 is a methodological artifact due to the top-100 selection constraint.
@@ 5.2.4 LLM-as-a-Judge’s limitation in reflecting human-perceived creativity
Failure under joint novelty–acceptability filtering. To further assess the ability of LLM-based judges to reflect human-perceived conceptual creativity, we jointly consider the dimensions of novelty and acceptability by explicitly incorporating multiple creativity-relevant criteria. Using multidimensional LLM evaluation prompts provided in Section G, we filter data points to retain only those for which both novelty-related and acceptability-related scores exceed predefined thresholds. The retained samples are treated as the predicted creative subset and the excluded set as the predicted uncreative subset. We then compare these predictions against human ground truth.
For Creative writing, we retain samples with LLM ratings of Logic $\geq 4$ ("Mostly logical and well-structured") and Originality $\geq 4$ (“Refreshingly different with clever elements”). For Research ideation, we retain samples with Presentation $\geq 3$, Soundness $\geq 3$, and Contribution $\geq 3$. As shown in Table 4, the resulting accuracies are 0.51 and 0.54 respectively—both close to random guessing. This indicates that even when jointly conditioning on multiple creativity-relevant dimensions, LLM-based judges do not robustly capture idea-level creativity.
Domain acc. F1 prec. rec. TP FP FN TN
Creative writing 0.51 0.04 1.00 0.02 3 0 147 150
Research ideation 0.54 0.48 0.55 0.42 218 180 295 333
Table 4: Classification metrics by domain. Accuracy (acc.), F1, precision (prec.), recall (rec.), and confusion-matrix counts are reported, with creative samples defined as the positive class (TP/FP/TN/FN).
@@
Figure 6: LLM evaluation results with F1, precision, recall, and confusion-matrix reported. (a) Problem-solving: LLMs tend to rate the solutions to be conventional; (b) Creative writing: LLMs tend to rate the plots to be non-original.
@@ 5.2.5 Improving LLM-as-a-Judge via chain-of-thought and rubric-based prompting
Chain-of-thought (CoT; Wei et al., 2022) assists LLMs in better adhering to evaluation instructions and consequently refining their final output scores. For example, in the third round of dialogue, the Claude agent revises (soundness, presentation, contribution, overall) evaluation scores from (3.5, 3, 3.5, 7) to (3, 3, 3, 6), with an explanation "…However, for a conference with only 30% acceptance rate, I must be stringent. The paper, while valuable, doesn’t appear to represent a breakthrough…, warranting a "Weak Accept" rating." This demonstrates the model’s ability to self-correct and adjust its assessment based on the provided context and constraints.
@@
5.2.6 Concept-level measurements: Semantic Novelty Score (SNS)
To explore potential approaches for evaluating conceptual creativity, we design an exploratory experiment, a two-stage pipeline that produces a Semantic Novelty Score (SNS), to investigate whether concept-level semantic distances can serve as a proxy for conceptual creativity. Specifically, we first employ gemini-2.5-flash as a semantic parser to extract the five most distinct core concepts ($|C|=5$) from each generated response. Second, we use the text-embedding-004 high-dimensional embedding space as a continuous proxy for a semantic graph, mapping the extracted concepts into this space. We then calculate the pairwise semantic distance, defined as 11- cosine similarity, across all unique concept pairs $P$. The SNS is defined as the average over these pairs:
$\mathrm{SNS}=\frac{1}{|P|}\sum_{(c_{i},c_{j})\in P}\left(1-\mathrm{sim}(v_{i},v_{j})\right).$
Table 5 presents evaluation results on subsets of 30 creative and 30 uncreative samples from each of the three domains. The inability of the SNS to improve discrimination across domains underscores the challenges faced by automatic metrics in capturing idea-level novelty, even when semantic novelty is explicitly targeted.
Domain Truth label SNS
Mean std.
Creative writing human expert 0.447 0.030
LLM-generated 0.421 0.084
Problem-solving unconventional 0.402 0.036
conventional 0.402 0.035
Research ideation accepted 0.420 0.049
rejected 0.435 0.037
Table 5: Semantic Novelty Score (SNS) by domain and label. SNS does not consistently distinguish between creative and uncreative outputs across domains.
@@ 6 Conclusion
Future research can build on our findings by collecting high-quality, orthogonal datasets in underexplored domains such as musical composition, humor, or product design. These datasets would complement our current benchmarks and support the development of generalizable, multi-domain, conceptually aware creativity judges that capture signals beyond surface-level patterns. Once these datasets are available, we can fine-tune LLMs on conceptual rather than syntactical signals. We suggest the next step may be incorporating recent advances such as Barrault et al. (2024)’s Large Concept Model (LCM), which uses sentence-level embeddings as a proxy for concept-level semantics, and continuous chain‐of‐thought architectures (Hao et al., 2025), which enable extended reasoning over idea chains. We believe that explainability could be a linchpin for creativity metrics that align more closely with human expectations and serve as a crucial component of future creativity evaluators.
Acknowledgments
The authors thank the PlusLab members at UCLA and the anonymous reviewers for their valuable feedback and helpful discussions. This work was supported in part by the National Science Foundation CAREER award #2339766, an Amazon AGI foundation research award, and the National Science and Technology Council, Taiwan, under the Grant 114-2628-E-002-021-, and the Taiwan Centers of Excellence. Shao-Hua Sun was supported by the Yushan Fellow Program of the Ministry of Education, Taiwan.
@@ Appendix A Data curation for scientific ideation
Alongside the accept or reject status, we collect the presentation, soundness, contribution, and overall scores for each paper. Scientific contribution is commonly used to represent creativity in the context of research ideation (Zhao and Zhang, 2025). To keep the focus on the research framing and ideation components, we filter for the Introduction section of papers, select papers based on the average of their overall review scores, and utilize soundness and presentation scores as control variables to reduce the influence of other factors.
@@ Appendix B CI implementation
For the results in Table 2, we compute CI with $L$-uniqueness in the range $5\leq L\leq 7$, following Appendix B.3 of Lu et al. (2024c). We observe the same saturation behavior for $L\geq 7$ (values $\approx$ 1), and therefore adopt the same setting. We also consider the near-verbatim match method using Word Mover’s Distance. However, this approach has computational complexity $O(|d|^{2}|w|)$ for each document $d$ and n-gram $w$, making it impractical for our large, diverse datasets. Moreover, (Lu et al., 2024c) reported similar human-vs-LLM gaps whether using verbatim or verbatim+semantic matching. Since parts of the original infrastructure (e.g., Elasticsearch WIMBD indices) of the “near-match CI” concept in Lu et al. (2024c) have been deprecated, we implemented a version using Llama-3 8B Instruct embeddings. Specifically, we apply text embeddings to the same index used for CI in our main experiments (v4_dolmasample_olmo), and evaluate this variant on balanced subsets of 30 creative and 30 uncreative samples from each of the three domains. The results in Table 6 closely mirror our CI findings: the creative and uncreative distributions overlap substantially, indicating that even the embedding-based evaluation is not discriminative as a metric for creativity. In addition, we observe the same sensitivity to the L-range, with CI scores increasing as $L$ grows, as reported in Section 5.2.2.
Domain Truth label L5 L6 L7 L5–6 L5–7
Creative Writing human expert 0.896 0.929 0.962 0.913 0.929
LLM-generated 0.878 0.921 0.963 0.900 0.921
Problem Solving unconventional 0.900 0.957 0.981 0.928 0.946
conventional 0.899 0.958 0.984 0.928 0.947
Research Ideation accepted 0.897 0.939 0.973 0.918 0.936
rejected 0.905 0.948 0.976 0.926 0.943
Table 6: Embedding-based CI performance by domain and truth label. Scores increase with larger $L$ but overlap substantially between labels across domains.
@@ Appendix C Label verification for creative writing
In the dataset from prior work (Tian et al., 2024a), turning points (markers of a story’s pacing) were shown to have highly significant differences at tp4 (mean_diff = 3.59-3.59, $p=8.7\times 10^{-8}$) and tp5 (mean_diff = 4.19-4.19, $p=5.4\times 10^{-8}$), indicating that LLM narratives reach these late-stage beats much earlier than human ones. Similarly, the distribution of story arcs (general plotlines such as “Man in Hole” or “Rags to Riches”) differs dramatically ($p=1.9\times 10^{-11}$). To further support our assumption, we have run some additional pairwise human annotations.
@@
Appendix D Human annotation instructions
We present samples from each domain together with detailed instructions for the human annotation tasks.
@@ Appendix D Human annotation instructions
Research ideation: "Below are some paper introductions. Select the one with higher contribution, bringing more new scientific breakthroughs and innovation to the research community."
@@ Appendix E Metrics yield misaligned performance across the same concept
[…]
@@ F.1 Prompt selection
Roleplaying with constraints vs. multi-aspect rubric-based prompt: Simply instructing the model does not result in more discriminative judgments. For example, in scientific ideation, contribution scores under constraints are indistinguishable from those produced with score-based novelty prompts. In contrast, multi-aspect rubric with orthogonal dimensions better isolates confounding factors such as presentation and soundness, and provides judgments more specifically targeted at scientific creativity, as discussed in Section 5.2.5.
@@ F.4 Potential data contamination
Group Truth label Human scores Claude 3 Haiku GPT-o1 GPT-4o GPT-o3-mini Claude Sonnet 3.7
@@
Table 9: Performance of LLM judges with different cutoffs. Comparison of human and models’ evaluation across groups (red = groups misaligned with truth labels; green = groups aligned with truth labels). Although Claude Sonnet 3.7 achieves the strongest alignment, this may be due to potential training data contamination rather than a genuine ability to evaluate conceptual novelty.
@@ Appendix G LLM evaluation prompts
Domain - Creative writing ”’Given a movie plot: “‘{MOVIE PLOT}“‘ Analyze the plot in terms of the following four aspects: Background Setup, Logic, Development, and Originality. For each aspect, consider the following questions: Background Setup: How common or unique is the background setting? Logic: Is the timeline clear and are the cause-and-effect relationships logical? Development: Are the twists well-developed? Is the buildup to the climax sufficient and engaging? Originality: How predictable is the plot? How original are the events and developments? For each aspect, discuss the strengths and weaknesses of the plot. Provide a thoughtful and critical analysis.”’, ”’Given the movie plot and your analyzation across four aspects, which The criteria for each aspect are as follows: A. Background Setup 1. Very common or clichéd background, lacks uniqueness. 2. Somewhat common, but with a few unique elements. 3. Generally unique, with some familiar elements. 4. Mostly unique background and setup. 5. Exceptionally unique background setup. B. Logic 1: Poor logic, many inconsistencies or plot holes. 2: Somewhat confusing or unrealistic transitions. 3: Generally coherent with minor issues. 4: Mostly logical and well-structured. 5: Strong internal logic and smooth cause-effect flow. C. Development 1: Underdeveloped or abrupt changes. 2: Weak buildup, limited depth. 3: Average development, a few effective moments. 4: Good buildup with engaging twists. 5: Excellent pacing, buildup, and well-executed climax. D. Originality 1: The plot is highly predictable or derivative. 2: Mostly familiar with few surprises. 3: Balanced mix of familiar and fresh ideas. 4: Refreshingly different with clever elements 5: Exceptionally creative and surprising or inspiring throughout. Ensure that each aspect is orthogonal to the others—for example, the uniqueness of background setup should not affect the evaluation of plot originality and development. You are a rigorous reviewer. Provide the scores in the following format: [[Background Setup: X]], [[Logic: X]], [[Development: X]], [[Originality: X]].”’
@@
Domain - Problem-solving ”’Given the problem: “‘{PROBLEM}“‘ and the solution: “‘{SOLUTION}“‘, analyze the following three aspects, taking into account the corresponding dimensions: 1. Problem Difficulty: Is the problem conceptually complex, or does it involve multiple steps or layers of reasoning? Are there hidden constraints or non-obvious elements that make the problem harder? 2. Unconventionality of the Solution: Does the solution follow a standard method, or does it use a surprising or creative strategy? Would this solution surprise someone familiar with typical methods for this problem? Are the tools, concepts, or constraints used in unexpected ways — beyond their typical or intended function? 3. Efficiency of the Solution: Does it reach the goal using the fewest steps or resources possible? Could the solution be improved or simplified without losing correctness? Carefully summarize and provide a thorough assessment of each aspect.”’, ”’Given the problem, solution, and your analyzation across three aspects, provide a score for each aspect accordingly, which The criteria for each aspect are as follows: A. Problem Difficulty 1: Very Easy (trivial or commonly seen problems; requires little reasoning) 2: Easy (straightforward with minimal abstraction or complexity) 3: Moderate (requires some thought or mild creativity; manageable) 4: Challenging (requires significant thought, creative insight, or multiple reasoning steps) 5: Very Challenging (highly complex or abstract; involves deep insight or several layers of reasoning) B. Unconventionality 1: Very Conventional (the most expected approach; standard method anyone would think of first) 2: Conventional (follows a known method that is not surprising, but slightly less obvious) 3: Moderately Unconventional (mix of familiar and unexpected methods or framings) 4: Unconventional (surprising to an expert; includes a creative or less typical approach) 5: Highly Unconventional (reframes the problem or uses tools and constraints in a deeply creative, unexpected way) C. Efficiency 1: Very Inefficient (excessive or redundant steps; far from optimal) 2: Inefficient (works, but could clearly be improved with fewer steps or resources) 3: Moderately Efficient (reasonable efficiency; could still be optimized) 4: Efficient (solid balance of steps and resources; minimal waste) 5: Very Efficient (achieves the goal in the fewest possible steps or resources with no clear room for improvement) Ensure that each aspect is orthogonal to the others—for example, Unconventionality should not be affected by the Efficiency or the Problem Difficulty. Strictly follow the criteria and provide the scores in the following format: [[Problem Difficulty: X]], [[Unconventionality: X]], [[Efficiency: X]].”’
@@
Domain - Research ideation ”’ You are a reviewer for a top conference, the acceptance rate for this conference is only around 20-30%. Given the introduction of the paper: “‘{PAPER CONTENT}“‘, carefully summarize and provide a thorough assessment of the strengths and weaknesses of the paper, touching on each of the following dimensions: 1. Originality: Are the tasks or methods new? Is the work a novel combination of well-known techniques? Is it clear how this work differs from previous contributions? 2. Quality: Is the submission technically sound? Are claims well supported (e.g., by theoretical analysis or experimental results)? Are the methods used appropriate? Is this a complete piece of work or work in progress? Are the authors careful and honest about evaluating both the strengths and weaknesses of their work? 3. Clarity: Is the submission clearly written? Is it well organized? Does it adequately inform the reader? 4. Significance: Are the results important? Are others (researchers or practitioners) likely to use the ideas or build on them? Does the submission address a difficult task in a better way than previous work? Does it advance the state of the art in a demonstrable way? Does it provide unique data, unique conclusions about existing data, or a unique theoretical or experimental approach? 5. Limitations: Have the authors adequately addressed the limitations and potential negative societal impact of their work? If not, please include constructive suggestions for improvement.”’, ”’Given the introduction of the paper and the strengths and weaknesses you summarized, analyze the four aspects of the paper repectively: Soundness, Presentation, Contribution, and Overall Assessment. The criteria for each aspect are as follows: A. Soundness: the soundness of the technical claims, experimental and research methodology and on whether the central claims of the paper are adequately supported with evidence. 4 excellent; 3.5; 3 good; 2.5; 2 fair; 1.5; 1 poor. B. Presentation: the quality of the presentation. This should take into account the writing style and clarity, as well as contextualization relative to prior work. 4 excellent; 3 good; 2 fair; 1 poor. C. Contribution: the quality of the overall contribution this paper makes to the research area being studied. Are the questions being asked important? Does the paper bring a significant originality of ideas and/or execution? Are the results valuable to share with the broader research community. 4 excellent; 3 good; 2 fair; 1 poor. D. Overall: 10: Award quality: Technically flawless paper with groundbreaking impact on one or more areas, with exceptionally strong evaluation, reproducibility, and resources, and no unaddressed ethical considerations. 9: Very Strong Accept: Technically flawless paper with groundbreaking impact on at least one area and excellent impact on multiple areas, with flawless evaluation, resources, and reproducibility, and no unaddressed ethical considerations. 8: Strong Accept: Technically strong paper with, with novel ideas, excellent impact on at least one area or high-to-excellent impact on multiple areas, with excellent evaluation, resources, and reproducibility, and no unaddressed ethical considerations. 7: Accept: Technically solid paper, with high impact on at least one sub-area or moderate-to-high impact on more than one area, with good-to-excellent evaluation, resources, reproducibility, and no unaddressed ethical considerations. 6: Weak Accept: Technically solid, moderate-to-high impact paper, with no major concerns with respect to evaluation, resources, reproducibility, ethical considerations. 5: Borderline accept: Technically solid paper where reasons to accept outweigh reasons to reject, e.g., limited evaluation. Please use sparingly. 4: Borderline reject: Technically solid paper where reasons to reject, e.g., limited evaluation, outweigh reasons to accept, e.g., good evaluation. Please use sparingly. 3: Reject: For instance, a paper with technical flaws, weak evaluation, inadequate reproducibility and incompletely addressed ethical considerations. 2: Strong Reject: For instance, a paper with major technical flaws, and/or poor evaluation, limited impact, poor reproducibility and mostly unaddressed ethical considerations. 1: Very Strong Reject: For instance, a paper with trivial results or unaddressed ethical considerations.”’, ”’Given the introduction of the paper, the strengths and weaknesses you summarized, and your analysis of the four aspects, provide a score for each aspect accordingly. Ensure that each aspect is orthogonal to the others—for example, Presentation should not affect the evaluation of Soundness or Contribution. You are a rigorous reviewer, and the conference has an acceptance rate of no more than 30%. Provide the scores in the following format:[[Soundness_llm: X]], [[Presentation_llm: X]], [[Contribution_llm: X]], [[Overall_llm: X]].”’